OpenStudy (anonymous):

Find f() if f(n) = 3n. 1/3 1 3

OpenStudy (anonymous):

No solutions exist

OpenStudy (anonymous):

So it would have to be 3

OpenStudy (anonymous):

are you looking for f(3) = 3*3 = 9

OpenStudy (anonymous):

do each one of them indvidiually to solve.

OpenStudy (anonymous):

Find f(1/3) if f(n) = 3n.

OpenStudy (anonymous):

i ment to put that

OpenStudy (anonymous):

plug in n for 1/3 and solve

OpenStudy (anonymous):

what did you get for the answer David?